OpenStudy (anonymous):

The population in the South in 2000 was 100,236,820. In 2010, the population in the South had increased by 14.3% from 2000. About many people lived in the South in 2010?

OpenStudy (anonymous):

let the population in 2000 be x. x=100236820 in 2010 the new population is= x+x*14.3%=x(1+14.3/100)=x*1.143 now find out your answer

OpenStudy (anonymous):

100,237,106?

OpenStudy (anonymous):

100,236,963

OpenStudy (anonymous):

carry out the multiplication. the answer is different than you gave. use calculator

OpenStudy (anonymous):

100,236,963?

OpenStudy (anonymous):

no, x=100236820, answer=x*1.143

OpenStudy (anonymous):

100,237,963

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:/ i dont understand?

OpenStudy (anonymous):

what am i doing wrong?

OpenStudy (anonymous):

the multiplication result is different. 100236820*1.143= do it please

OpenStudy (anonymous):

oh multiply it i see hold up

OpenStudy (anonymous):

114,570,685.26?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok thank you so much

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

I need one more if you can help me out?

OpenStudy (anonymous):

please it is my pleasure

OpenStudy (anonymous):

The population in the Northeast was 53,594,378 people in 2000 and 55,317,240 people in 2010. By about what percent did the population change? Did the population grow or shrink?

OpenStudy (anonymous):

how would i solve that?

OpenStudy (anonymous):

let the population in 2000 be x and in 2010 be y. x=53,594,378 and y=55,317,240 here x>y. so population decreases. the amount of shrink is=[(x-y)/x]*100% please do it.

OpenStudy (anonymous):

so subtract x from y?

OpenStudy (anonymous):

sorry, I have missed. y>x. and the amount will be [(y-x)/x]*100%

OpenStudy (anonymous):

so y subtract x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1722862?

OpenStudy (anonymous):

sorry for no commas

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

this is increment. the result is =[(y-x)/x]*100%

OpenStudy (anonymous):

i dont understand?

OpenStudy (anonymous):

you want me to divide that by x?

OpenStudy (anonymous):

yes devide it by x. then multiply with 100. the result will come

OpenStudy (anonymous):

wait give me the layout like on how to do it?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

1722862*100/x is the answer

OpenStudy (anonymous):

17,228,62/53,594,378 x 100?

OpenStudy (anonymous):

i got 3.214632?

OpenStudy (anonymous):

yes 3.214632% grow

OpenStudy (anonymous):

Thank you so much because of you i might pass algebra ...tomorrow is the end of my semester

OpenStudy (anonymous):

ok, you will do well.

OpenStudy (anonymous):

Thank you again:)

OpenStudy (anonymous):

Im a fan now of yours

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

So the population shrunk right?

OpenStudy (anonymous):

no it will grow as y>x i.e. population 2010> population 2000

OpenStudy (anonymous):

Oh yeah lol i see now thanks again

OpenStudy (anonymous):

I did another one i just want to know if i was right

OpenStudy (anonymous):

The 2010 census reported about 308.7 million people in the United States. The 2000 census reported a population of about 281.4 million. By about what percent did the population change? Did the population grow or shrink?

OpenStudy (anonymous):

my answer was 2730%

OpenStudy (anonymous):

Can you give me an answer to one thing i need it to pass algebra by midnight

OpenStudy (anonymous):

A compact car will drive three miles east, five miles west, and then 13 miles east again. The car gets 40 miles per gallon, and there is 1 4 of a gallon of gas left in the tank. Compute the relative distance and the total distance. Then decide if there is enough fuel for the trip.

OpenStudy (anonymous):

oh my god you're my savior

OpenStudy (anonymous):

2730% is wrong. do as 100(y-x)/x

OpenStudy (anonymous):

its ok i already turned it in

OpenStudy (anonymous):

with x=281.4 million and y=308.7 million

OpenStudy (anonymous):

but can you help me with this man if i dont turn it in before midnight im going to fail algebra i do online school so its different

OpenStudy (anonymous):

A compact car will drive three miles east, five miles west, and then 13 miles east again. The car gets 40 miles per gallon, and there is 1 4 of a gallon of gas left in the tank. Compute the relative distance and the total distance. Then decide if there is enough fuel for the trip.

OpenStudy (anonymous):

i do not understand 1 4 of a gallon means what?

OpenStudy (anonymous):

the rough diagram is, if - represents 1 mile and right side is east --- towards east ----- towards west ------------- towards east so total is 3+5+13 miles. relative for 3 mile east and 5 mile west is= 3-5 east=-2 east then relative is (13)+(-2)=11 east

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

so is that the answer? i dont understand?

OpenStudy (anonymous):

I really appreciate the help it means alot

OpenStudy (anonymous):

i have given only total distance and relative distance. the car travels only total distance. 1 gallon can cover 40 mile. 1/4 gallon can cover 10 mile. so 1/4 gallon is present in the car at final position. so the if it wants to come back directly initial position it has to cover relative distance i.e. 11 mile. but it has the fuel for 10 mile. so if you understand it, say the trip is possible or not

OpenStudy (anonymous):

its not possible

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

Here is another one

OpenStudy (anonymous):

A van will drive ten miles north, 15 miles south, and then five miles north again. The van gets 33 miles per gallon, and there is one gallon of gas left in the tank. Compute the relative distance and the total distance. Then decide if there is enough fuel for the trip.

OpenStudy (anonymous):

the total is just add i.e. 10+15+5 miles. the relative is considering direction 10north+15south+5north =10north-15north+5north now decide if the trip is possible or not seeing relative distance only. your previous understanding was right.

OpenStudy (anonymous):

It is possible

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you yes just one more will you help me one last time?:)

OpenStudy (anonymous):

Last one

OpenStudy (anonymous):

A sedan will drive ten miles north, 15 miles east, 13 miles south, 15 miles west, and 22 miles north. The sedan gets 20 miles per gallon, and there are five gallons of gas left in the tank. Compute the relative distance and the total distance. Then decide if there is enough fuel for the trip.

OpenStudy (anonymous):

total distance=10+15+13+15+22 miles relative=10north+15east+13south+15west+22north =10north+15east-13north-15east+22north =19north 1 gallon cover 20 miles. 5 gallon cover 5*20=100 miles. now you say.

OpenStudy (anonymous):

Yes its possible?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Im so thankful you helped me is there a way i can repay you or anything? i wanna keep in touch with you do you play like xbox or anything like that lol

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol what do you like to do?

OpenStudy (anonymous):

ahh i need help with this

OpenStudy (aravindg):

14.3% increased means increased number is (14.3/100 )*2000 So population in 2010 is 2000+(14.3/100 )*2000 Solve it.